OpenStudy (anonymous):

verify the identity: (1-2 sin^2 θ)/(sin θ cos θ)=cot θ-tan θ

OpenStudy (anonymous):

cotx-tanx= cosx/sinx - sinx/cosx. Which becomes cos^2x-sin^2x/sinxcosx. cos^2x-sin^2x=1-2sin^2x. So the right side now looks like 1-2sin^2x/sinxcosx. And you're done

OpenStudy (anonymous):

let x=theta

OpenStudy (anonymous):

so 1-2sin^2x/sinxcosx is the identity?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

\[\cot \Theta = \frac{ \cos \Theta }{ \sin \Theta }\] \[\tan \Theta = \frac{ \sin \Theta }{ \cos \Theta }\] \[\frac{ \cos ^{2}\Theta - \sin ^{2} \Theta }{ \cos \Theta \sin \Theta }\] alright so far?

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yeah thanks!

OpenStudy (anonymous):

Good! :) You're welcome

OpenStudy (anonymous):

\[\cos ^{2} \Theta = 1- \sin ^{2} \Theta\] Insert that into the cos pat and wala\[\frac{ 1 - \sin ^{2} \Theta }{ \cos \Theta \sin \Theta}\]